# Sharp thresholds for spanning regular subgraphs

### v2

@@ 6.4. Reconstruction from a smoothed fragment
$\sum_{\ell\geq\lceil n^{0.3}\rceil+1}\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\sum_{c,\ell,x}\left[\left(\frac{n^{0.2}}{c}\right)^{c}e^{-C\ell}+\left(\frac{n^{0.55}}{(c+\sigma)^{2}}\right)^{c+\sigma}\right]$
$\leq\sum_{\ell\geq n^{0.3}}\left(e^{-\ell}+\sum_{x}\left(\sum_{c\leq n^{0.25}}n^{0.2c}e^{-C\ell}+\sum_{c>n^{0.25}}n^{-0.05c}\right)\right)$
@@ 6.6. Day 2: getting fragments of size $o(\log n)$
For $\ell\in\mathbb{N}$ and $H\in\mathcal{H}^{(i)}$, let $\Pi^{H}_{\ell}:=\mathbb{P}(H\cap\mathbf{H}\text{ has }\ell\text{ edges})$. We let $f=f_{i}$ and $\mathcal{B}$ be the set of all graphs on $n$ vertices with more than $f_{i+1}$ edges. By Lemma 4.1, it is sufficient to show (40) with $\lceil n^{0.3}\rceil$ replaced by $f_{i+1}$. For non-negative integers $c$ and $x$, we denote, as above, the probability that $\mathbf{H}\cap H\text{ has }\ell\text{ edges, }x\text{ non-isolated vertices, and }c\text{ components}$ by $p^{H}(\ell,x,c)$. Here, letting $\sigma:=2x-3c-\ell$ and $\ell_{\chi}:=\ell_{0}+5\chi$, we have that $n^{c/6}<{\ell_{\chi}\choose c+\sigma}$. Therefore, due to Claims 3.26.7, we get
@@
Note that we have not yet specified the value of $w$. The only restriction that we have is $1\ll w\ll\log n$. We, therefore, may further assume that $\frac{\log n}{\log\log\log n}\ll w\ll\log n$. In particular, it implies the following refinement of the inequality (42):
@@
$e^{(\ln\ln\ln n)^{3}}<\sqrt{\ln n}<\frac{f_{i}^{1/3}}{2(\ln\ln\ln n)^{2}}\quad\text{ and }\quad f_{i+1}+1\geq(f_{i}-1)^{2/3}+1\geq f_{i}^{2/3},$
@@ 6.7. Day 3: covering a fragment
On the other hand, recalling that $\mathcal{J}^{H}_{\ell,x,c}$ is defined in Section 6.5, due to Claim 6.7, letting $\ell_{\chi}:=\ell_{0}+5\chi$, we get
$\mathrm{Var}Y^{(R+1)}\leq|\mathcal{H}^{(R+1)}|\sum_{\ell,x,c}\max_{H\in\mathcal{H}^{(R+1)}}|\mathcal{J}^{H}_{\ell,x,c}|64^{\ell}(n-x+c)!\frac{\ell^{c+\sigma}_{\chi}}{(\ln n)^{c+\sigma}}(p^{\prime})^{2\lceil(\ln n)^{0.3}\rceil-\ell}-(\mathbb{E}Y^{(R+1)})^{2},$
where $\sigma=2x-3c-\ell$. This implies
$\frac{\mathrm{Var}Y^{(R+1)}}{(\mathbb{E}Y^{(R+1)})^{2}}\leq\sum_{\ell,x,c}\max_{H\in\mathcal{H}^{(R+1)}}\frac{|\mathcal{J}^{H}_{\ell,x,c}|\cdot 64^{\ell}(n-x+c)!\ell^{c+\sigma}_{\chi}}{(1-o(1))n!p^{\prime\ell}(\ln n)^{c+\sigma}}-1.$
@@
$\frac{\mathrm{Var}Y^{(R+1)}}{(\mathbb{E}Y^{(R+1)})^{2}}$ $\leq\sum_{\ell\geq 0,c,x}\frac{{\lceil(\ln n)^{0.3}\rceil\choose c}2^{12\ell}(2C\sqrt{n})^{c+\sigma}}{n^{\ell/2+c/2+\sigma/2}(\varepsilon\cdot n^{-1/2})^{\ell}(\ln n)^{c+\sigma}}-1$
$\leq\sum_{\ell\geq 1,c,x}\frac{(e\lceil(\ln n)^{0.3}\rceil)^{c}\left(2^{13}C/\varepsilon\right)^{\ell}}{(\ln n)^{c+\sigma}}\leq\sum_{\ell\geq 1,c,x}(\ln n)^{-(c+\sigma)/2}$
@@
completing the proof due to Chebyshev’s inequality.
Now, let $\mathbf{W}^{\prime}\sim G(n,m^{\prime})$ , where $m^{\prime}=\lceil p^{\prime}N\rceil$. Then whp here exists $H\in\mathcal{H}^{(R+1)}$ such that $H\subset\mathbf{W}^{\prime}$. We conclude that whp $\mathbf{W}\cup\mathbf{W}^{(1)}\cup\ldots\cup\mathbf{W}^{(R)}\cup\mathbf{W}^{\prime}$ contains the second power of a Hamilton cycle. Then, we apply  [17, Corollary 1.16] in the opposite direction and get that whp $\mathbf{G}\sim G(n,(1+10\varepsilon)\sqrt{e/n})$ contains the second power of a Hamilton cycle, as well.
@@ 6.8. Proof of Claim 34
Let $\ell\geq\ell_{0}+1$ be an integer.
We first note that the square of a path with $\ell$ edges has $\frac{\ell+3}{2}$ vertices. Therefore, a disjoint union of $c$ squares on $\ell$ edges has $\frac{\ell+3c}{2}$ vertices. We then fix integers $c\in[\ell]$ and $x\in\left[\frac{\ell}{2}+\frac{3}{2}c,\ell+c\right]$. For a subgraph $H\subset F\in\mathcal{F}_{n}$ with $x$ vertices, $\ell$ edges, and $c$ components,
$\sigma:=2x-\ell-3c$
@@
Let $p(\ell,x,c)$ be the probability that $(\mathbf{F}\cap F)\cup(D_{1}\cup\ldots\cup D_{\chi})$ is a graph on $x+4\chi$ vertices with $\ell+5\chi$ edges, and $c+\chi$ connected components (as usual, we think about graphs as about sets of their edges, so there are no isolated vertices in $\mathbf{F}\cap F$). The following claim immediately implies that, for sufficiently large $A_{1},A_{2}>0$,
$p(\ell,x,c)\leq\frac{{n\choose c}{x\choose c+\chi}\left(\frac{2x}{\chi}\right)^{\chi}e^{A_{1}c+A_{2}\sigma}\left(\max_{o\leq 8\sigma}{x\choose o}\right)^{2}(n-x-4\chi+c)!}{|\mathcal{F}_{n}(\overrightarrow{D})|}.$ (45)
@@
There exist $A^{1}_{\alpha},A^{2}_{\alpha}>0$ such that the number of subgraphs in $F$ without isolated vertices with $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components is
$\alpha(\ell,x,c)\leq{n\choose c}{x\choose c+\chi}\left(\frac{2x}{\chi}\right)^{\chi}e^{A^{1}_{\alpha}c+A^{2}_{\alpha}\sigma}\max_{o\leq 8\sigma}{x\choose o}.$ (46)
There exist $A^{1}_{\beta},A^{2}_{\beta}>0$ such that, given $H\subset F$ with $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components, the number of ways to extend $H$ to a graph from $\mathcal{F}_{n}(\overrightarrow{D})$ is at most
@@
Fix such $x_{i}$ and $\ell_{i}$, $i\in[c+\chi]$. For $i\in[c+\chi]$, set $\sigma_{i}=2x_{i}-\ell_{i}-3$.
We first choose closed subgraphs of $F$ that correspond to $\sigma_{i}=0$ one by one. The number of ways to choose the first closed subgraph is at most $2n$ (the heaviest case is $x_{i}=2$ that correspond to an edge). In the same way, if some set of $\tilde{n}$ vertices is already included in the subgraph under construction, then the next closed component can be chosen in at most $2(n-\tilde{n})$ ways. On the other hand, if the $i$-th component contains one of the diamonds (i.e. $i\geq c+1$), then the number of ways to choose it is at most $x_{i}$.
@@
    choose the number of free vertices $o_{i}\leq 6+2\sigma_{i}\leq 8\sigma_{i}\leq 8^{\sigma_{i}}$;
@@
    choose a set $\mathcal{O}\in{[x_{i}]\choose o_{i}}$ that identifies the labels of free vertices in the $i$-th component;
@@
    if $i\leq c$ choose a vertex $w$ out of the set of remaining $n-\tilde{n}$ vertices and, if $i\geq c+1$, choose $w$ such that $10\leq\pi_{F}(u_{i-c})-\pi_{F}(w)\leq x_{i}-$ 1, where $u_{i-c}$ is the minimum vertex of the diamond $D_{i-c}$ that this component has to contain; activate $w$ — we treat this vertex as the $\pi_{F}$-minimum vertex in the component under construction;
@@
We get that, if $i\leq c$, then the number of ways to choose the component is at most $(n-\tilde{n})8^{\sigma_{i}}\max\limits_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}$, and at most $x_{i}8^{\sigma_{i}}\max\limits_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}$, otherwise.
@@
$c!{n\choose c}2^{c}(x/\chi)^{\chi}\prod_{i=1}^{c+\chi}8^{\sigma_{i}}\max_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}\leq c!{n\choose c}(x/\chi)^{\chi}2^{c}2^{35\sigma}\max_{o\leq 8\sigma}{x\choose o}.$
@@
$\alpha(\ell,x,c)$ ($\leq{n\choose c}{x\choose c+\chi}{\sigma+c+\chi\choose c+\chi}\left(\frac{x}{\chi}\right)^{\chi}2^{c+35\sigma}\max_{o\leq 8\sigma}{x\choose o}$
($\leq{n\choose c}{x\choose c+\chi}\left(\frac{2x}{\chi}\right)^{\chi}2^{2c+36\sigma}\max_{o\leq 8\sigma}{x\choose o}$
@@
Let us now fix $H\subset F$ with $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components. Let us bound the number of ways to extend $H$ to an $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D})$. We construct such an extension in the following way.
We forget the labels of the vertices from $V(H)\setminus V(D_{1}\cup\ldots\cup D_{\chi})$ and assume that the desired $F^{\prime}\in\mathcal{F}_{n}$ is the second power of the cycle $(12\ldots n)$. Let $\mathcal{Z}$ be the union of the set of the $c$ connected components of $H$ that do not contain diamonds with the set of remaining $n-x-4\chi$ isolated vertices. We should compute the number of ways to embed the elements of $Z$ in $F^{\prime}$ disjointly.
Let $z_{1},\ldots,z_{n-x-4\chi+c}$ be an ordering of $Z$ (there are $(n-x-4\chi+c)!$ ways to order the elements of $Z$). We embed sequentially each $z_{i}$ in $F^{\prime}$ in a way such that all vertices of $z_{i}$ are bigger than the minimum vertices of $z_{1},\ldots,z_{i-1}$. At every step $i=1,\ldots,n-x-4\chi+c$, consider the minimum vertex $\kappa_{i}$ of $F^{\prime}$ such that none of the embedded elements of $Z$ in $F^{\prime}$ contain this vertex. If $z_{i}$ is a single vertex, then we assign $\kappa_{i}$ with $z_{i}$ and proceed with the next step. Otherwise, we distinguish between the following cases.
@@
Second, let $z_{i}$ be not closed with $o$ free vertices. Choose a set $\mathcal{O}$ from (${[|V(z_{i})|]\choose o}$ that will identify the labels of free vertices in the embedding of $z_{i}$ in $F^{\prime}$. Activate $\kappa_{i}$. At every step $j\geq 1$, choose the $\pi_{F^{\prime}}$-minimum active vertex $v_{j}$ in $F^{\prime}$ and
  • if $j\in\mathcal{O}$, then add to the image of $z_{i}$ under construction some neighbours of $v_{j}$ (in at most $2^{3}$ ways), deactivate $v_{j}$, and activate all its neighbours,
  • if $j\notin\mathcal{O}$, then add all the neighbours of $v_{j}$, deactivate the vertex, and activate all its neighbours.
The image is constructed. However, we have not yet mapped the vertices of $z_{i}$ to the vertices of the image. Clearly, $z_{i}$ is union of a ‘path’ $P$ of inclusion-maximum 2-connected graphs joined by (usual) paths with ‘fringe’ trees, growing from $P$. There are two ways to orient $P$, unless $P$ is a cycle. Every rooted ‘fringe’ tree $T$ has at most $2^{|V(T)|-1}$ automorphism since $T$ is a subtree of a perfect 3-ary tree. Moreover, the number of free non-root vertices in $T$ is at least $|V(T)|/2$. In total, there are at most $2\cdot 2^{2o-1}=2^{2o}$ automorphisms of $z_{i}$ and this bound holds true even when $P$ is a cycle. We get that the number of ways to construct the image of $z_{i}$ is at most (${|V(z_{i})|\choose o}2^{5o}$.
It is easy to see that, for a component of $H$ that contains some $D_{i}$ and has $o$ free vertices, there are at most $6^{o}$ ways to embed it in $F^{\prime}$. We conclude that there are at most
$(n-x-4\chi+c)!6^{c}\prod_{i=1}^{c}{x_{i}\choose o_{i}}2^{5o_{i}}\prod_{i=c+1}^{c+\chi}6^{o_{i}}\leq(n-x-4\chi+c)!6^{c}2^{40\sigma}\max_{o\leq 8\sigma}{x\choose o}$
@@
Recalling that $|\mathcal{F}_{n}(\overrightarrow{D})|=(n-4\chi)!$ and $x-c=\ell/2+c/2+\sigma/2$, we get
@@
$\left(\frac{2x}{\chi}\right)^{\chi}=e^{\chi\ln(2x/\chi)}=e^{\chi\ln(x/\chi)+\chi\ln 2}=e^{o(x)}=e^{o(\ell)}.$
We first choose $0<\varepsilon^{\prime}\ll\delta$ small enough. From (45) and the bound (${n\choose c}\leq\left(\frac{en}{c}\right)^{c}$, we get

### v3

@@ 6.4. Reconstruction from a smoothed fragment
$\sum_{\ell\geq\lceil n^{0.3}\rceil+1}\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\sum_{c,\ell,x}\left[\left(\frac{n^{0.2}}{c}\right)^{c}e^{-C\ell/2}+\left(\frac{n^{0.55}}{(c+\sigma)^{2}}\right)^{c+\sigma}\right]$
$\leq\sum_{\ell\geq n^{0.3}}\left(e^{-\ell}+\sum_{x}\left(\sum_{c\leq n^{0.25}}n^{0.2c}e^{-C\ell/2}+\sum_{c>n^{0.25}}n^{-0.05c}\right)\right)$
@@ 6.6. Day 2: getting fragments of size $o(\log n)$
For $\ell\in\mathbb{N}$ and $H\in\mathcal{H}^{(i)}$, let $\Pi^{H}_{\ell}:=\mathbb{P}(H\cap\mathbf{H}\text{ has }\ell\text{ edges})$. We let $f=f_{i}$ and $\mathcal{B}$ be the set of all graphs on $n$ vertices with more than $f_{i+1}$ edges. By Lemma 4.1, it is sufficient to show (45) with $\lceil n^{0.3}\rceil$ replaced by $f_{i+1}$. For non-negative integers $c$ and $x$, we denote, as above, the probability that $\mathbf{H}\cap H\text{ has }\ell\text{ edges, }x\text{ non-isolated vertices, and }c\text{ components}$ by $p^{H}(\ell,x,c)$. Here, letting $\sigma:=2x-3c-\ell$ and $\ell_{\chi}:=\ell_{0}+5\chi$, we have that $c+\sigma=2(x-c)\ell\leq\ell<\ell_{\chi}/2$ and also
$n^{c/6}{\ell_{\chi}\choose\sigma}=n^{c/6}\frac{(c+\sigma)!(\ell_{\chi}-c-\sigma)!}{\sigma!(\ell_{\chi}-\sigma)!}{\ell_{\chi}\choose c+\sigma}$ $<\left(n^{1/6}\cdot\frac{c+\sigma}{\ell_{\chi}-c-\sigma}\right)^{c}{\ell_{\chi}\choose c+\sigma}$
$<\left(\frac{n^{1/6}\ell}{\ell_{\chi}-\ell}\right)^{c}{\ell_{\chi}\choose c+\sigma}<{\ell_{\chi}\choose c+\sigma}.$
Therefore, due to Claims 3.26.7, we get
@@
Note that we have not yet specified the value of $w$. The only restriction that we currently have is $1\ll w\ll\log n$. We, therefore, may further assume that $\frac{\log n}{\log\log\log n}\ll w\ll\log n$. In particular, it implies the following refinement of the inequality (48):
@@
$e^{(\ln\ln\ln n)^{3}}<(\ln n)^{0.1}<\frac{f_{i}^{1/3}}{2(\ln\ln\ln n)^{2}}\quad\text{ and }\quad f_{i+1}+1\geq(f_{i}-1)^{2/3}+1\geq f_{i}^{2/3},$
@@ 6.7. Day 3: covering a fragment
On the other hand, recalling that $\mathcal{J}^{H}_{\ell,x,c}$ is defined in Section 6.5, due to the second part of Claim 6.7 (recall that in the beginning of Section 6.5 we chose the fragments satisfying the bound (44) deterministically), letting $\ell_{\chi}:=\ell_{0}+5\chi$ and $\sigma:=2x-3c-\ell$, we get
$\mathrm{Var}Y^{(R+1)}\leq|\mathcal{H}^{(R+1)}|\sum_{\ell,x,c}\max_{H\in\mathcal{H}^{(R+1)}}|\mathcal{J}^{H}_{\ell,x,c}|64^{\ell}(n-x+c)!\frac{(2\ell_{\chi})^{c+\sigma}}{(\ln n)^{c+\sigma}}(p^{\prime})^{2\lceil(\ln n)^{0.3}\rceil-\ell}-(\mathbb{E}Y^{(R+1)})^{2}.$
This implies
$\frac{\mathrm{Var}Y^{(R+1)}}{(\mathbb{E}Y^{(R+1)})^{2}}\leq\sum_{\ell,x,c}\max_{H\in\mathcal{H}^{(R+1)}}\frac{|\mathcal{J}^{H}_{\ell,x,c}|\cdot 64^{\ell}(n-x+c)!(2\ell_{\chi})^{c+\sigma}}{(1-o(1))n!p^{\prime\ell}(\ln n)^{c+\sigma}}-1.$
@@
$\frac{\mathrm{Var}Y^{(R+1)}}{(\mathbb{E}Y^{(R+1)})^{2}}$ $\leq\sum_{\ell\geq 0,c,x}\frac{{\lceil(\ln n)^{0.3}\rceil\choose c}2^{12\ell}(4C\sqrt{n})^{c+\sigma}}{n^{\ell/2+c/2+\sigma/2}(\varepsilon\cdot n^{-1/2})^{\ell}(\ln n)^{c+\sigma}}-1$
$\leq\sum_{\ell\geq 1,c,x}\frac{(e\lceil(\ln n)^{0.3}\rceil)^{c}\left(2^{14}C/\varepsilon\right)^{\ell}}{(\ln n)^{c+\sigma}}\leq\sum_{\ell\geq 1,c,x}(\ln n)^{-(c+\sigma)/2}$
@@
completing the proof that an $H\subset\mathbf{G}^{\prime}$ whp exists, due to Chebyshev’s inequality.
Now, let $\mathbf{W}^{\prime}\sim G(n,m^{\prime})$ , where $m^{\prime}=\lceil p^{\prime}N\rceil$. Then whp there exists $H\in\mathcal{H}^{(R+1)}$ such that $H\subset\mathbf{W}^{\prime}$. We conclude that whp $\mathbf{W}\cup\mathbf{W}^{(1)}\cup\ldots\cup\mathbf{W}^{(R)}\cup\mathbf{W}^{\prime}$ contains the square of a Hamilton cycle. Then, we apply  [18, Corollary 1.16] in the opposite direction and get that whp $\mathbf{G}\sim G(n,(1+10\varepsilon)\sqrt{e/n})$ contains the square of a Hamilton cycle, as well.
@@ 6.8. Proof of Claim 39
Let $\ell\geq\ell_{0}+1$ be an integer. We first note that the square of a path with $\ell$ edges has $\frac{\ell+3}{2}$ vertices. Therefore, a disjoint union of $c$ squares on $\ell$ edges has $\frac{\ell+3c}{2}$ vertices. We then fix integers $c\in[\ell]$ and $x\in\left[\frac{\ell}{2}+\frac{3}{2}c,\ell+c\right]$. For a subgraph $H\subset F\in\mathcal{F}_{n}$ with $x$ vertices, $\ell$ edges, and $c$ components,
$\sigma:=2x-\ell-3c$ (53)
@@
Now, let $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and let $\mathbf{F}$ be a uniformly random element of $\mathcal{F}^{*}_{n}(\overrightarrow{D})$. Let $p(\ell,x,c)$ be the probability that $(\mathbf{F}\cap F)\cup(D_{1}\cup\ldots\cup D_{\chi})$ is a graph on $x+4\chi$ vertices with $\ell+5\chi$ edges, and $c+\chi$ connected components (as usual, we think about graphs as about sets of their edges, so there are no isolated vertices in $(\mathbf{F}\cap F)\cup(D_{1}\cup\ldots\cup D_{\chi})$). The following claim immediately implies that, for sufficiently large $A_{1},A_{2}>0$,
$p(\ell,x,c)\leq\frac{{n\choose c}{x\choose c+\chi}\left(\frac{4x+2\chi}{\chi}\right)^{\chi}e^{A_{1}c+A_{2}\sigma}\left(\max_{o\leq 8\sigma}{x\choose o}\right)^{2}(n-x-4\chi+c)!}{|\mathcal{F}_{n}(\overrightarrow{D})|}.$ (54)
@@
There exist $A^{1}_{\alpha},A^{2}_{\alpha}>0$ such that the number of subgraphs $H\subset F$ such that $H\cup(D_{1}\cup\ldots\cup D_{\chi})$ is a graph without isolated vertices with $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components is
$\alpha(\ell,x,c)\leq{n\choose c}{x\choose c+\chi}\left(\frac{4x+2\chi}{\chi}\right)^{\chi}e^{A^{1}_{\alpha}c+A^{2}_{\alpha}\sigma}\max_{o\leq 8\sigma}{x\choose o}.$ (55)
There exist $A^{1}_{\beta},A^{2}_{\beta}>0$ such that, given $H\subset F$ such that $H\cup(D_{1}\cup\ldots\cup D_{\chi})$ has $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components, the number of ways to extend $H$ to a graph from $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ is at most
@@
Fix such $x_{i}$ and $\ell_{i}$, $i\in[c+\chi]$. For $i\in[c]$, set $\sigma_{i}=2x_{i}-\ell_{i}-3$, while, for $i\in[c+\chi]\setminus[c]$, set $\sigma_{i}=2x_{i}-\ell_{i}$.
Let us compute the number of ways to choose connected vertex-disjoint subgraphs $R_{1},\ldots,R_{c+\chi}$ from $F$ such that, for every $i\in[c]$, $R_{i}$ has $\ell_{i}$ edges and $x_{i}$ vertices, and, for every $i\in[c+\chi]\setminus[c]$, $R_{i}\cup D_{i}$ is a connected graph with $\ell_{i}+5$ edges and $x_{i}+4$ vertices. We first choose closed subgraphs of $F$ that correspond to $\sigma_{i}=0$ one by one. The number of ways to choose the first closed subgraph is at most $2n$ (the heaviest case is $x_{i}=2$ that corresponds to an edge). In the same way, if some set of $\tilde{n}$ vertices is already included in the subgraph under construction, then the next closed component can be chosen in at most $2(n-\tilde{n})$ ways. On the other hand, if the $i$-th component contains one of the diamonds (i.e. $i\geq c+1$), then the number of ways to choose it is at most $x_{i}+1$.
@@
    choose the number of free vertices $o_{i}\leq 6+2\sigma_{i}\leq 8\sigma_{i}\leq 8^{\sigma_{i}}$ (recall that $\sigma_{i}\geq 1$);
@@
    choose a set $\mathcal{O}\in{[x_{i}]\choose o_{i}}$ that comprises the labels of free vertices in the $i$-th component;
@@
    if $i\leq c$, choose a vertex $w$ outside of the set of the remaining $n-\tilde{n}$ vertices and, if $i\geq c+1$, choose $w$ such that $0\leq\pi_{F}(u_{i-c})-\pi_{F}(w)\leq 2x_{i}$, where $u_{i-c}$ is the minimum vertex of the diamond $D_{i-c}$ that this component has to contain; activate $w$ — we treat this vertex as the $\pi_{F}$-minimum vertex in the component under construction;
@@
We get that, if $i\leq c$, then the number of ways to choose the component is at most $(n-\tilde{n})8^{\sigma_{i}}\max\limits_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}$, and at most $(2x_{i}+1)8^{\sigma_{i}}\max\limits_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}$, otherwise.
@@
$c!{n\choose c}2^{c}\left(\frac{2x+\chi}{\chi}\right)^{\chi}\prod_{i=1}^{c+\chi}8^{\sigma_{i}}\max_{o_{i}\leq 8\sigma_{i}}{x_{i}\choose o_{i}}2^{4o_{i}}\leq c!{n\choose c}\left(\frac{2x+\chi}{\chi}\right)^{\chi}2^{c+35\sigma}\max_{o\leq 8\sigma}{x\choose o}.$
@@
$\alpha(\ell,x,c)$ ($\leq{n\choose c}{x\choose c+\chi}{\sigma+c+\chi\choose c+\chi}\left(\frac{2x+\chi}{\chi}\right)^{\chi}2^{c+35\sigma}\max_{o\leq 8\sigma}{x\choose o}$
($\leq{n\choose c}{x\choose c+\chi}\left(\frac{4x+2\chi}{\chi}\right)^{\chi}2^{2c+36\sigma}\max_{o\leq 8\sigma}{x\choose o}$
@@
Let us now fix $H\subset F$ such that $H\cup(D_{1}\cup\ldots\cup D_{\chi})$ has $x+4\chi$ vertices, $\ell+5\chi$ edges, and $c+\chi$ components. Let us bound the number of ways to extend $H$ to an $F^{\prime}\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$. We construct such an extension in the following way.
We forget the labels of the vertices from $V(H)\setminus V(D_{1}\cup\ldots\cup D_{\chi})$ and assume, without loss of generality, that the desired $F^{\prime}\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ is such that $F^{\prime}\cup D_{1}\cup\ldots\cup D_{\chi}$ is the square of the cycle $(12\ldots n)$. Let $\mathcal{Z}$ be the union of the set of the $c$ connected components of $H$ that do not contain diamonds with the set of remaining $n-x-4\chi$ isolated vertices. Let $\mathcal{Z}^{*}$ be the union of components of $H$ that contain some $D_{i}$. Then the desired number of extensions is exactly the number of ways to embed the elements of $\mathcal{Z}\sqcup\mathcal{Z}^{*}$ in $F^{\prime}$ disjointly, where in each component in $\mathcal{Z}^{*}$ only the vertices of diamonds are labelled.
Let $z_{1},\ldots,z_{n-x-4\chi+c}$ be an ordering of $\mathcal{Z}$ (there are $(n-x-4\chi+c)!$ ways to order the elements of $\mathcal{Z}$). We embed sequentially each $z_{i}$ in $F^{\prime}$ in a way such that all vertices of $z_{i}$ are bigger than the minimum vertices of $z_{1},\ldots,z_{i-1}$. At every step $i=1,\ldots,n-x-4\chi+c$, consider the minimum vertex $\kappa_{i}$ of $F^{\prime}$ such that none of the embedded elements of $\mathcal{Z}$ in $F^{\prime}$ contain this vertex. If $z_{i}$ is a single vertex, then we assign $\kappa_{i}$ with $z_{i}$ and proceed to the next step. Otherwise, we distinguish between the following cases.
@@
Second, let $z_{i}$ be not closed with $o_{i}$ free vertices. Choose a set $\mathcal{O}_{i}$ from (${[|V(z_{i})|]\choose o_{i}}$ that comprises the labels of free vertices in the embedding of $z_{i}$ in $F^{\prime}$ (we forget the isomorphism class of $z_{i}$ at this point and estimate the number of ways to expose a connected subgraph of $F^{\prime}$ with $|V(z_{i})|$ vertices, $o_{i}$ free vertices, starting from $\kappa_{i}$). Activate $\kappa_{i}$. At every step $j\geq 1$, choose the $\pi_{F^{\prime}}$-minimum active vertex $v_{j}$ in $F^{\prime}$ and
  • if $j\in\mathcal{O}_{i}$, then add to the image of $z_{i}$ under construction some neighbours of $v_{j}$ (in at most $2^{3}$ ways), deactivate $v_{j}$, and activate all its added neighbours,
  • if $j\notin\mathcal{O}_{i}$, then add all the neighbours of $v_{j}$, that have not been added yet, deactivate the vertex, and activate all its added neighbours.
The image is constructed. However, we have not yet mapped the vertices of $z_{i}$ to the vertices of the image. Clearly, $z_{i}$ is union of a ‘path’ $P$ of inclusion-maximum 2-connected graphs, joined by (usual) paths, with ‘fringe’ trees, growing from $P$. There are at most 4 automorphisms of $P$, unless $P$ is a cycle. Since the closed 1-neighbourhood of every vertex in a square of a cycle does not have an independent set of size 3, every non-root vertex of a rooted ‘fringe’ tree $T$ has degree at most 3 — so it is free. In particular, $T$ has at most $2^{|V(T)|-1}$ automorphisms since it is a subtree of a perfect binary tree. We conclude that, in total, there are at most $4\cdot 2^{o_{i}}=2^{o_{i}+2}\leq 2^{2o_{i}}$ automorphisms of $z_{i}$ and this bound holds true even when $P$ is a cycle. We get that the number of ways to construct the image of $z_{i}$ is at most (${|V(z_{i})|\choose o_{i}}2^{5o_{i}}$.
We finally switch to embedding components from $\mathcal{Z}^{*}$, which is an easier task, since they are “rooted” in advance. Indeed, it is easy to see that, for a component of $H$ that contains some $D_{i}$ and has $o_{i}$ free vertices, there are at most $6^{o_{i}}$ ways to embed it in $F^{\prime}$. Recalling the bound $o_{i}\leq 6+2\sigma_{i}\leq 8\sigma_{i}\mathbbm{1}_{\sigma_{i}\geq 1}$, we conclude that there are at most
$(n-x-4\chi+c)!6^{c}\times\prod_{i\in[c]:\,\sigma_{i}\geq 1}{x_{i}\choose o_{i}}2^{5o_{i}}\times\prod_{i\in[c+\chi]\setminus[c]:\,\sigma_{i}\geq 1}6^{o_{i}}\leq(n-x-4\chi+c)!6^{c}2^{40\sigma}\max_{o\leq 8\sigma}{x\choose o}$
@@
Let us now finish the proof of Claim 39. Recalling that $|\mathcal{F}_{n}(\overrightarrow{D})|=(n-4\chi)!$ and that (53) implies
$x-c=((x+4\chi)-(c+\chi))-3\chi=((\ell+5\chi)/2+(c+\chi)/2+\sigma/2)-3\chi=\ell/2+c/2+\sigma/2,$
we get
@@
$\left(\frac{4x+2\chi}{\chi}\right)^{\chi}=e^{\chi\ln\left(\frac{x}{\chi}\cdot\frac{4x+2\chi}{x}\right)}=e^{\chi\ln(x/\chi)+\chi\ln(4+o(1))}=e^{o(x)}=e^{o(\ell)}.$
We are now ready to derive (39). Let us choose $0<\varepsilon^{\prime}\ll\delta$ small enough. From (54) and the bound (${n\choose c}\leq\left(\frac{en}{c}\right)^{c}$, we get